OpenStudy (anonymous):

Summation Evaluate these summations. a) N(E)k=1 (k-k^3) b) N-1(E)k=1 (k^2/N^2) Note that (E) is the summation symbol. Not sure how to do these. Help would be appreciated

OpenStudy (anonymous):

hey m not sure i understand..r the terms on the right the expression for the kth term??

OpenStudy (anonymous):

do we have to evaluate the sum of the series whose term is k-k^3????clarify and i cn help

OpenStudy (anonymous):

sorry it's all squashed together, didn't notice! there are two questions: a) N (E) k=1 (k-k^3) b) N-1 (E) k=1 (k^2/N^2) where N = natural numbers and (E) is the summation symbol The N and N-1 for the questions are on top of the summation symbol and the k =1 are below the symbol. The (k-k^3) and (k^2/N^2) are on the right sides of the summation symbol.

OpenStudy (anonymous):

u can use the equation tab to write it in mathematical format

OpenStudy (anonymous):

ugghh it keeps squashing together - but a) and b) are separate questions

OpenStudy (anonymous):

N(N+1)/2 + [N(N+1)/2]^2

OpenStudy (anonymous):

thats a)

OpenStudy (anonymous):

as for b) (N-1)(N)(2N-1)/6N ^2

OpenStudy (anonymous):

\[\sum_{k=1}^{N}(k-k^3)\]

OpenStudy (anonymous):

yes uzma that's the correct format for a)

OpenStudy (anonymous):

and b?????

OpenStudy (anonymous):

\[\sum_{k}^{N-1}(k^2/N^2)\]

OpenStudy (anonymous):

i gave d answers

OpenStudy (anonymous):

Yes that's the correct one for b) him1618, thanks for your help :) and uzma too :)

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

did u understand the solution?

OpenStudy (anonymous):

Yes, thank you :)

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

wht tym is it there at ur place?

OpenStudy (anonymous):

9:02pm

OpenStudy (anonymous):

where r u?

OpenStudy (anonymous):

new zealand

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

N(E)K=1(K-K^3) (E)1/1-K^2=1/N N=1-K^2 K=1 N=0

OpenStudy (anonymous):

what sort of followup is this kihaga???

OpenStudy (anonymous):

Yes! that is Honorata kihaga